OpenStudy (anonymous):

The first part of a lecture on volcanoes lasts 1/2 hour. The next two parts each last 1/4 hour. There are breaks of 5 minutes between each part. If you must leave after 55 minutes, how much of the lecture will you miss, explain?

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (anonymous):

@StudyGurl14 Could you help

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (anonymous):

@StudyGurl14 this is math now, haha

OpenStudy (anonymous):

hhhheeeeeeeeeeeeeeeeeeeeeelllllllllllllllllooooooooooooooo

OpenStudy (studygurl14):

\(\large\frac{5}{60}\rightarrow\frac{1}{12}\) \(\large\frac{55}{60}\rightarrow\frac{11}{12}\) \(\Large (\frac{1}{2}+\frac{1}{4}+\frac{1}{4}+2(\frac{1}{12}))-\frac{11}{12}=\tt the~answer\)

OpenStudy (anonymous):

so -11/12 or 11/12

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

What are you talking about?

OpenStudy (anonymous):

is it -11/12 or 11/12

OpenStudy (studygurl14):

Next step: \(\Large 1+\frac{2}{12}-\frac{11}{12}=\tt the~answer\) ...because \(\large\frac{1}{2}+\frac{1}{4}+\frac{1}{4}=1\)

OpenStudy (anonymous):

so what is the answer?

OpenStudy (studygurl14):

Next Step: \(\Large 1-\frac{9}{12}=\tt the~answer\) ...because \(\large\frac{2}{12}-\frac{11}{12}=-\frac{9}{12}\)

OpenStudy (studygurl14):

And since \(\large 1=\frac{12}{12}\) \(\Large\frac{12}{12}-\frac{9}{12}\rightarrow\frac{3}{12}\rightarrow\frac{1}{4}=\tt{the~answer}\)

OpenStudy (studygurl14):

Can I have a medal now?

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

Just kidding